# Exploring How Generative MLLMs Perceive More Than CLIP with the Same Vision Encoder

### v2

@@ 1 Introduction
In this work, we observe that the query-relevant visual information could still be preserved by CLIP vision encoder, but a better strategy is required to extract it: As shown in Figure 1, LLaVA-1.5-7B (Liu et al., 2024) with the same pretrained vision encoder, surpasses CLIP-ViT-L/14-336px by a large margin on many challenging visual reasoning benchmarks. Particularly, on spatial reasoning benchmark What’sUp, while CLIP’s pair accuracy is lower than random chance (25%), LLaVA-1.5 achieves beyond 50% on all four subsets (Table 2.1). More evidence on various benchmarks and other Generative MLLMs showing this phenomenon is presented in Section 2. These results indicate that these Generative MLLMs extract and utilize query-relevant information more effectively from the same CLIP vision encoder. Notably, the vision encoder remains unchanged throughout training, ensuring a fair comparison.
@@ 2.1 Task Setup
Models. Our main comparison is between CLIP-ViT-L/14-336px (Radford et al., 2021) and Generative MLLMs that use its pretrained vision encoder and keep the weights frozen during training: LLaVA-1.5-7B (Liu et al., 2024), along with Phi-3-V-3.8B and LLaMA-3-V-8B (Rasheed et al., 2024). In these MLLMs, the patch tokens from the CLIP vision encoder first pass through a two-layer MLP connector and are then used as input tokens for a generative language model which yields the token probability determining the model response. We also include results of CLIP-ViT-L/14-224px and SigLIP-ViT-L/16-384px (Zhai et al., 2023) for reference since they are of interest and widely used (Tong et al., 2024a).
@@ 3.1 Training Data
Table 4: The two-way individual accuracy and pair accuracy results of CLIP-ViT-L/14-336px and SigLIP-ViT-L/16-384px focusing on the Left/Right subsets of What’sUp after finetuning on LLaVA-1.5’s training data with or without hard negative captions. After direct finetuning, the accuracies are still quite low.
First, we hypothesize that Generative MLLMs’ visual information extraction ability benefits from training data. To check the effect of data, we use LLaVA-1.5’s training data to finetune CLIP and SigLIP. We convert the datasets to the image-caption format (Details are deferred to Appendix B.2). By default, we freeze the vision encoder during finetuning for strict ablation. Considering that CLIP training relies on negative samples beyond data quality (Robinson et al., 2020; Kalantidis et al., 2020), we also construct hard negative captions by switching the related phrases to their opposite (e.g., replacing “on the left” with “on the right”). In this setting, the training objective follows NegCLIP (Yuksekgonul et al., 2023).
Results are shown in Table 4. Finetuning on LLaVA-1.5’s training data does not help these two models, even with hard negatives. Still, their accuracy is around random chance. We also try to unlock the SigLIP vision encoder during finetuning, which does not increase the performance either (See results in Appendix B.3). We experiment with XVLM (Zeng et al., 2021) and observe similar results in Appendix B.4. This finding aligns with the previous failure on finetuning them on a much larger, preposition-focused subset of LAION (Kamath et al., 2023b), indicating that data alone does not lead to stronger extraction ability.
@@ 3.2 Token Usage
We first perform an ablation study on LLaVA-1.5: We change the input of its language model to use only the [CLS] token, train this “[CLS]-LLaVA-1.5” model from scratch (pretraining + finetuning) using LoRA (Hu et al., 2021), and observe that its spatial reasoning performance is significantly worse than our reproduced LLaVA-1.5-LoRA in Table 5. This proves the importance of patch tokens to fine-grained visual reasoning: Detailed information of images reside in these patch tokens.
@@ 3.3 Language Model
Are pretrained LLMs the missing piece to effectively extracting visual information? We perform further experiments on finetuning CLIP with patch tokens and RoPE on LLaVA-1.5 training data but replacing the original CLIP text encoder with a stronger one provided by LLM2CLIP (Huang et al., 2024). This text encoder is converted from Llama-3-8B-Instruct (Dubey et al., 2024) by contrastive finetuning and is shown to bring performance boost to state-of-the-art CLIP models on benchmarks such as MS COCO (Lin et al., 2014). We keep this text encoder and the CLIP vision encoder frozen during our finetuning. The results are shown in the fifth row of Table 6, where we also attach the results of the original LLM2CLIP checkpoint of CLIP-ViT-L/14-336px for reference [^1] LLM2CLIP is not a fair comparison as its implementation unfreezes the CLIP vision encoder during finetuning.. We find that a stronger text encoder does not suffice to effectively extract more information towards solving the task.
@@ 3.4 Alignment Architecture, Training Objective, and Prompt
We bypass this obstacle in comparison by converting a Generative MLLM to a CLIP-like contrastive VLM. On LLaVA-1.5-7B, we use the converting method proposed by VLM2Vec (Jiang et al., 2024): Specifically, we take the last layer vector representation of the last output token of LLaVA-1.5-7B as the output embedding. In this way, we get the encoder for image+question(prompt) and pure text simultaneously since LLaVA-1.5 allows using one or zero images in the input. Following the original paper, we use the prompt templates: “Represent the given image with the following question: {Question}” while encoding the image if there is a question in the sample; “Find the text that can answer the given query: {Question}” when there is no image; and no additional prompt for encoding image+question of LCS-558K and the captions. Then, we finetune this encoder using contrastive loss and LoRA (Hu et al., 2021) on LLaVA-1.5’s training data with the CLIP vision encoder frozen. Surprisingly, they exhibit strong performance without using a large batch size (256) in the second row of Table 3.3 (LLaVA-1.5-7B-VLM2Vec-LoRA). The questions used for evaluation is listed in Appendix A.4. This proves that text generation+autoregressive loss is not the only solution to fine-grained visual reasoning.
@@ Acknowledgments
We thank Amita Kamath for the discussion about the What’sUp benchmark. We appreciate the comments and advice from Rui Xin and Scott Geng on our drafts. PWK is supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Innovation under the AI Visiting Professorship Programme (award number AIVP-2024-001). SSD acknowledges the support of NSF DMS 2134106, NSF CCF 2212261, NSF IIS 2143493, NSF IIS 2229881, and Sloan Fellowship.
@@ A.1 Benchmark Information
The What’sUp benchmark (Kamath et al., 2023b) contains 820 images of pairs of household objects captured by the authors, 408 in Subset A and 412 in Subset B. For every object pair, all prepositions are present in the benchmark, and thus the images and captions are balanced, avoiding the bias in real-world images (e.g., a cup is usually on the table, not under the table). We corrected the mislabeled images in the GitHub Issues and reevaluated the pretrained VLMs. For CLIP, SigLIP, and XVLM’s evaluation, we refer to the official code provided by the What’sUp benchmark’s authors in https://github.com/amitakamath/whatsup_vlms. We offer an example in Figure 3 to demonstrate how pair accuracy and individual accuracy are computed on benchmarks with paired images like What’sUp.
@@ A.2 Model Weight Information
We use public pretrained weights of LLaVA-1.5-7B (https://huggingface.co/llava-hf/llava-1.5-7b-hf) under the Meta LLaMA License Agreement and the weights of Phi-3-V-3.8B in https://huggingface.co/MBZUAI/LLaVA-Phi-3-mini-4k-instruct and LLaMA-3-V-8B in https://huggingface.co/MBZUAI/LLaVA-Meta-Llama-3-8B-Instruct provided by (Ranasinghe et al., 2024) under MIT License since they are trained with vision encoder frozen. For contrastive VLMs, we use OpenAI’s pretrained CLIP-ViT-L/14-224px and CLIP-ViT-L/14-336px model under MIT License, and SigLIP-ViT-L/16-384px pretrained on the WebLI dataset (Chen et al., 2022) under Apache 2.0 License provided in the OpenCLIP repository. In Table 5, LLaVA-1.5-7B-LoRA is reproduced. In Table 6, the checkpoint for LLM2CLIP is from https://huggingface.co/microsoft/LLM2CLIP-Llama-3-8B-Instruct-CC-Finetuned under Apache 2.0 License. We also use the text encoder and the adapter of this checkpoint in our experiments of using a stronger text encoder. In Table 3.3, the LLaVA-1.5-7B-VLM2Vec-LoRA is trained by ourselves with the vision encoder frozen using the VLM2Vec method (Jiang et al., 2024).
@@ A.3 Comparison between VQAScore and Response-Based Evaluation
Table 8: Question formats for different subsets for evaluating LLaVA-1.5-7B.
@@
Table 10: Results of CLIP-ViT-L/14-336px and Generative MLLMs on four subsets in What’sUp. The individual accuracy and pair accuracy are in percentage points.
@@ B.1 Hyperparameters
Our code for training standard CLIP and SigLIP is based on https://github.com/mlfoundations/open_clip (Ilharco et al., 2021). We finetune CLIP and SigLIP models for five epochs with a learning rate of 5e-6 on the combination of converted LCS-558K plus converted DataMix-665K. We use 50 steps of warmup and AdamW optimizer with a cosine-annealing learning rate schedule. The batch size is 512, and we train the models on 4 GPUs. The training time is less than one day.
@@
For the experiments in Table 6, we start from an implementation of PACL (Mukhoti et al., 2023) in https://github.com/NMS05/Patch-Aligned-Contrastive-Learning. Since we only need to train the vision embedder and text embedder, we apply a larger batch size (4096) and train for 10 epochs on 8 GPUs on the combination of converted LCS-558K plus converted DataMix-665K. We use 0.1 as the fixed temperature, 1e-4 as the learning rate, and Adam as the optimizer. The training time is less than one day for all experiments. We adopt the data augmentation implemented in the codebase, except the RandomHorizontalFlip, which discourages models from learning about Left/Right spatial relations. When processing the captions, we follow the technique used in the codebase, which uses the original caption randomly with a probability of 0.5, and a random template+(a noun in the caption) otherwise. The templates are: “a picture of {}.”, “itap of {}.”, “a photograph of {}.”, “this picture contains {}.”, “a good photo of {}.”. For experiments with a stronger text encoder, we do not apply this technique on DataMix-665K.

### v3

@@ 1 Introduction
In this work, we observe that the query-relevant visual information could still be preserved by CLIP vision encoder, but a better strategy is required to extract it: As shown in Figure 1, LLaVA-1.5-7B (Liu et al., 2024) with the same pretrained vision encoder, surpasses CLIP-ViT-L/14-336px by a large margin on many challenging visual reasoning benchmarks. Particularly, on spatial reasoning benchmark What’sUp, while CLIP’s pair accuracy is lower than random chance (25%), LLaVA-1.5 achieves beyond 50% on all four subsets (Table 2.1). More evidence of other Generative MLLMs on various benchmarks showing this phenomenon is presented in Section 2. These results indicate that these Generative MLLMs extract and utilize query-relevant information more effectively from the same CLIP vision encoder. Notably, the vision encoder remains unchanged throughout training, ensuring a fair comparison.
@@ 2.1 Task Setup
Models. Our main comparison is between CLIP-ViT-L/14-336px (Radford et al., 2021) and Generative MLLMs that use its pretrained vision encoder and keep the weights frozen during training: LLaVA-1.5-7B (Liu et al., 2024), along with Phi-3-V-3.8B and LLaMA-3-V-8B (Rasheed et al., 2024). In these MLLMs, the patch tokens from the CLIP vision encoder first pass through a two-layer MLP connector and are then used as input tokens for a generative language model which yields the token probability determining the model response. We also include results of CLIP-ViT-L/14-224px, SigLIP-ViT-L/16-384px (Zhai et al., 2023), and EVA01-ViT-g-14 (Sun et al., 2023) for reference since they are of interest and widely used (Tong et al., 2024a).
@@ 3.1 Training Data
Table 4: The two-way individual accuracy and pair accuracy results of CLIP-ViT-L/14-336px, SigLIP-ViT-L/16-384px, and EVA01-ViT-g-14 focusing on the Left/Right subsets of What’sUp after finetuning on LLaVA-1.5’s training data with or without hard negative captions. After direct finetuning, the accuracies are still quite low.
First, we hypothesize that Generative MLLMs’ visual information extraction ability benefits from training data. To check the effect of data, we use LLaVA-1.5’s training data to finetune CLIP, SigLIP, and EVA-CLIP. We convert the datasets to the image-caption format (Details are deferred to Appendix B.2). By default, we freeze the vision encoder during finetuning for strict ablation. Considering that contrastive learning relies on negative samples beyond data quality (Robinson et al., 2020; Kalantidis et al., 2020), we also construct hard negative captions by switching the related phrases to their opposite (e.g., replacing “on the left” with “on the right”). In this setting, the training objective follows NegCLIP (Yuksekgonul et al., 2023).
Results are shown in Table 4. Finetuning on LLaVA-1.5’s training data does not help these models, even with hard negatives. Still, their accuracy is around random chance. We also try to unlock the SigLIP vision encoder during finetuning, which does not increase the performance either (See results in Appendix B.3). We experiment with XVLM (Zeng et al., 2021) and observe similar results in Appendix B.4. This finding aligns with the previous failure on finetuning them on a much larger, preposition-focused subset of LAION (Kamath et al., 2023b), indicating that data alone does not lead to stronger extraction ability.
@@ 3.2 Token Usage
We first perform an ablation study on LLaVA-1.5: We change the input of its language model to use only the [CLS] token, train this “[CLS]-LLaVA-1.5” model from scratch (pretraining + finetuning) using LoRA (Hu et al., 2021), and observe that its spatial reasoning performance is significantly worse than our reproduced LLaVA-1.5-LoRA in Table 5. This proves the importance of patch tokens to fine-grained visual reasoning: Detailed information of images resides in these patch tokens.
@@ 3.3 Language Model
Are pretrained LLMs the missing piece to effectively extracting visual information? We perform further experiments on finetuning CLIP with patch tokens and RoPE on LLaVA-1.5 training data but replacing the original CLIP text encoder with a stronger one provided by LLM2CLIP (Huang et al., 2024). This text encoder is converted from Llama-3-8B-Instruct (Dubey et al., 2024) by contrastive finetuning and is shown to bring performance boost to state-of-the-art CLIP models on benchmarks such as MS COCO (Lin et al., 2014). We keep this text encoder and the CLIP vision encoder frozen during our finetuning. The results are shown in the fifth row of Table 6, where we also attach the results of the original LLM2CLIP checkpoint of CLIP-ViT-L/14-336px for reference [^1] The original LLM2CLIP is not a fair comparison as its implementation unfreezes the CLIP vision encoder during finetuning.. We find that a stronger text encoder does not suffice to effectively extract more information towards solving the task.
@@ 3.4 Alignment Architecture, Training Objective, and Prompt
We bypass this obstacle in comparison by converting a Generative MLLM to a CLIP-like contrastive VLM. On LLaVA-1.5-7B, we use the converting method proposed by VLM2Vec (Jiang et al., 2024): Specifically, we take the last layer vector representation of the last output token of LLaVA-1.5-7B as the output embedding. In this way, we get the encoder for image+question(prompt) and pure text simultaneously since LLaVA-1.5 allows using one or zero images in the input. Following the original paper, we use the prompt templates: “Represent the given image with the following question: {Question}” while encoding the image if there is a question in the sample; “Find the text that can answer the given query: {Question}” when there is no image; and no additional prompt for encoding im- age+question of LCS-558K and the captions. Then, we finetune this encoder using contrastive loss and LoRA (Hu et al., 2021) on LLaVA-1.5’s training data with the CLIP vision encoder frozen. Surprisingly, they exhibit strong performance without using a large batch size (256) in Table 3.3 (LLaVA-1.5-7B-VLM2Vec-LoRA). The question used for evaluation is listed in Appendix A.4. This proves that text generation+autoregressive loss is not the only solution to fine-grained visual reasoning.
@@ Acknowledgments
We thank Amita Kamath for the discussion about the What’sUp benchmark. We appreciate the comments and advice from Rui Xin and Scott Geng on our drafts. PWK is supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Innovation under the AI Visiting Professorship Programme (award number AIVP-2024-001) and by the AI2050 program at Schmidt Sciences. SSD acknowledges the support of NSF DMS 2134106, NSF CCF 2212261, NSF IIS 2143493, NSF IIS 2229881, Sloan Fellowship, and the AI2050 program at Schmidt Sciences.
@@ A.1 Benchmark Information
The What’sUp benchmark (Kamath et al., 2023b) contains 820 images of pairs of household objects captured by the authors, 408 in Subset A and 412 in Subset B. For every object pair, all prepositions are present in the benchmark, and thus the images and captions are balanced, avoiding the bias in real-world images (e.g., a cup is usually on the table, not under the table). We corrected the mislabeled images in the GitHub Issues and reevaluated the pretrained VLMs. For CLIP and XVLM’s evaluation, we refer to the official code provided by the What’sUp benchmark’s authors in https://github.com/amitakamath/whatsup_vlms. The evaluation of SigLIP and EVA-CLIP directly follows the evaluation of CLIP in the official code. We offer an example in Figure 3 to demonstrate how pair accuracy and individual accuracy are computed on benchmarks with paired images like What’sUp.
@@ A.2 Model Weight Information
We use public pretrained weights of LLaVA-1.5-7B (https://huggingface.co/llava-hf/llava-1.5-7b-hf) under the Meta LLaMA License Agreement and the weights of Phi-3-V-3.8B in https://huggingface.co/MBZUAI/LLaVA-Phi-3-mini-4k-instruct and LLaMA-3-V-8B in https://huggingface.co/MBZUAI/LLaVA-Meta-Llama-3-8B-Instruct provided by (Ranasinghe et al., 2024) under MIT License since they are trained with vision encoder frozen. For contrastive VLMs, we use OpenAI’s pretrained CLIP-ViT-L/14-224px and CLIP-ViT-L/14-336px model under MIT License, SigLIP-ViT-L/16-384px pretrained on the WebLI dataset (Chen et al., 2022) and EVA01-ViT-g-14 pretrained on the LAION400M-s11b-b41k dataset (Schuhmann et al., 2021) under Apache 2.0 License provided in the OpenCLIP repository. In Table 5, LLaVA-1.5-7B-LoRA is reproduced. In Table 6, the checkpoint for LLM2CLIP is from https://huggingface.co/microsoft/LLM2CLIP-Llama-3-8B-Instruct-CC-Finetuned under Apache 2.0 License. We also use the text encoder and the adapter of this checkpoint in our experiments of using a stronger text encoder. In Table 3.3, the LLaVA-1.5-7B-VLM2Vec-LoRA is trained by ourselves with the vision encoder frozen using the VLM2Vec method (Jiang et al., 2024).
@@ A.3 Comparison between VQAScore and Response-Based Evaluation
Table 8: Question formats for different subsets for LLaVA-1.5-7B.
@@
Table 10: Results of CLIP-ViT-L/14-336px and Generative MLLMs on four subsets in What’sUp using standard response-based evaluation. The individual accuracy and pair accuracy are in percentage points.
@@ A.4 Evaluating VLM2Vec
In addition, we show that the benefit of using a question in the prompt generalizes beyond What’sUp. Here, we perform the same comparison as in Table 3.3 on MMVP and MMVP-VLM. We use the original question of the benchmark in the prompt. For MMVP-VLM which does not have questions, we manually add an MMVP-like question to each test case without altering content or tuning the prompt. We attached these questions to the updated supplementary material. We use the same prompt format as What’sUp (“Represent the given image with the following question: {Question}” or “Represent the given image.” without any question). We observe similar results in Table A.4.
MMVP MMVP-VLM
CLIP-ViT-L/14-336px 14.0 20.7
[…]
w/o Question in Prompt 9.3 11.9
Random chance 25.0 25.0
Table 11: The pair accuracy of CLIP-ViT-L/14-336px and LLaVA-1.5-converted models in percentage points on MMVP and MMVP-VLM. LLaVA-1.5-7B-VLM2Vec-LoRA continues to outperform CLIP. When there is no question in the prompt, its performance degenerates to the standard CLIP.
@@ B.1 Hyperparameters
Our code for training standard CLIP, SigLIP, and EVA-CLIP is based on https://github.com/mlfoundations/open_clip (Ilharco et al., 2021). We finetune these models for five epochs with a learning rate of 5e-6 on the combination of converted LCS-558K plus converted DataMix-665K. We use 50 steps of warmup and AdamW optimizer with a cosine-annealing learning rate schedule. The batch size is 512, and we train the models on 4 GPUs. The training time is less than one day.
@@
For the experiments in Table 6, we start from an implementation of PACL (Mukhoti et al., 2023) in https://github.com/NMS05/Patch-Aligned-Contrastive-Learning. Since we only need to train the vision embedder and text embedder, we apply a larger batch size (4096) and train for 10 epochs on 8 GPUs on the combination of converted LCS-558K plus converted DataMix-665K. We use 0.1 as the fixed temperature, 1e-4 as the learning rate, and Adam as the optimizer. The training time is less than one day for all experiments. We adopt the data augmentation implemented in the codebase, except the RandomHorizontalFlip, which discourages models from learning about Left/Right spatial relations. When processing the captions, we follow the technique used in the codebase, which uses the original caption randomly with a probability of 0.5, and template+(a noun in the caption) otherwise. The templates are: “a picture of {}.”, “itap of {}.”, “a photograph of {}.”, “this picture contains {}.”, “a good photo of {}.”. For experiments with a stronger text encoder, we do not apply this technique on DataMix-665K.